OpenStudy (anonymous):

There is an equal number of boys and girls in a 9th grade class. The ratio of the average heights of boys to girls is 12 to 11. If the average height of the girls in the class is 66 inches, what is the average height, in inches, of the boys in the class?

OpenStudy (anonymous):

this question requires you to setup a proportion \[\frac{ 12 }{ 11}=\frac{ boys }{girls }\] now how you solve this problem is set the proportion equal to the girls height with x as the boys height\[\frac{ 12 }{ 11}=\frac{ x }{ 66 }\] then cross multiply and get \[12(66) = 11x\] and you can take it from here